When testing the effectiveness of a medicine, 57.3% of those taking the medicine noticed an improvement in their health. For a random sample of 392 individuals, find the mean, variance, and standard deviation for the number of individuals who will not notice an improvement in their health. Show your work to receive credit.
Answer #1

Solution :

Given that,

p = 0.573

q = 1 - p =1-0.573=0.427

n = 392

Mean = = n * p = 392*0.573=224.616

variance 2= n * p * q =  392*0.573*0.427=95.9110

Standard deviation = = n * p * q =  392*0.573*0.427=9.7934